OpenStudy (anonymous):

a train 100m long is moving at a speed of 50 km/h crosses a train 120m long coming from opposite direction on parallel tracks in 6 sec. the speed of the second train is

OpenStudy (anonymous):

Would love to help, but this question is lacking a question... :/

ganeshie8 (ganeshie8):

any method u familiar wid @master71 ?

ganeshie8 (ganeshie8):

@techhelper i think we need to find the speed of second train

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ah, alright. i think he was still going to give more information. :)

ganeshie8 (ganeshie8):

say, speed of second train = \(x\) km/h

ganeshie8 (ganeshie8):

relative speed = \(50 + x\) km / h

ganeshie8 (ganeshie8):

overlapping distance = 100 + 120 = 220 next, use the relation :- distance = speed x time

ganeshie8 (ganeshie8):

can u solve \(x\) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

what is the ans

ganeshie8 (ganeshie8):

distance = speed x time \(\frac{220}{1000} = (50+x) \times \frac{6}{60\times 60} \) solve \(x\)

OpenStudy (anonymous):

ok thanks

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

the ratio of the length and breadth of a plot is 3:2 if the bredth ia 40m less than the length , what is the perimeter of the plot